If the momentum of the NASA space shuttle as it leaves the atmosphere is 3.75 * 10^8 kg * m/s and its mass is 75000 kg, what is its speed?
Answer #1

here,

mass , m = 75000 kg

let the speed of shuttle be v

the momentum when it leaves the atmosphere , P = m * v

3.75 * 10^8 = 75000 * v

v = 5 * 10^3 m/s

the speed of space shuttle is 5 * 10^3 m/s